OpenStudy (anonymous):

Pls help me on this PLS Write an equation of the line that is parallel to the graph of y-3x=4 and has the same y-intercept as the graph of 4y+x=36.

OpenStudy (anonymous):

when you say parallel, therefor, their slope are the same. . (1) find the y intercept of the second equation 4y=-x+36 y= -1/4 x +9 y-intercept is 9 (2) find the slope of the first equation y=3x+4 m(slope)=3 therefor the, using the slope=3 and the point (0,b) which is (0,9) ,since the y-intercept is 9, therefor the equation parallel to the first equation and the same y-intercept of the second equation is y=3x +9